Problem 10.06 Sunland Incorporated management is considering investing in two alternative production systems. The systems are mutually exclusive, and the cost of the new equipment and the resulting cash flows are shown in the accompanying table. The firm uses a 7 percent discount rate for their production systems Year System 1 System 2 $13,600 13,600 13,600 13,600 $46,200 32,400 32,400 32,400 0 2 What are the payback periods for production systems 1 and 2? (Round answers to 2 decimal places, e.g. 15.25.) Payback period of System 1 is years and Payback period of System 2 is years If the systems are mutually exclusive and the firm always chooses projects with the lowest payback period, in which system should the firm invest? The firm should invest in Click if you would like to Show Work for this question
Answer #1
Payback period is the period by which firm is able to recover its initial investment from cash flow of the project
Computation of payback period
System 1 payback period = 13600/13600= 1.00 year
System 2 payback period = 1+(46200-32400)/32400          1.43 year
The firm should invest in system 1 as it has lower payback period compared to system 2